find the thirteenth term of the following sequence -6, -4, -2, 0....
Answer by flame8855(424)

the sequence has the following : a1= -6 , d= 2
i think you meant the 30th term
a(n)= a1+(n-1)d
a(13)= -6 +(30-1)2= 52